# Numerical simulations of density perturbation and gravitational wave production from cosmological first-order phase transition

### v1

@@ Abstract
We conducted three-dimensional lattice simulations to study the density perturbation and generation of gravitational waves (GWs) through first-order phase transition (FOPTs). Our findings reveal that for the case of $\beta/H=6$, density perturbations successfully collapse into PBHs. Moreover, smaller values of $\beta/H$ are more conducive to PBH formation, while the strength of the phase transition $\alpha$ has a relatively minor impact on PBH production. We find that for $\alpha>1$, the dominant contribution to the density perturbation comes from the delay of vacuum decay, whereas for $\alpha<1$, the bubble walls’ moving forward behavior becomes the primary source. Additionally, the power spectrum of density perturbations generated by the phase transition exhibits a slope of $k^{3}$ at small wavenumbers and $k^{-1.5}$ at large wavenumbers. Furthermore, we calculated the GW power spectra, and the simulation results show a slope of $k^{3}$ at small wavenumbers and $k^{-2}$ at large wavenumbers. Our numerical simulations confirm that slow PTs can produce PBHs and provide predictions for the GW power spectrum, offering crucial theoretical support for GW detection.
@@ I INTRODUCTION
First-order phase transition (FOPT) is one important source of stochastic gravitational wave (GW) background Hindmarsh et al. (2021). During the FOPT, the anisotropic stresses produced by the bubble collisions Cutting et al. (2018, 2021), sound waves Hindmarsh et al. (2015, 2017, 2014); Cutting et al. (2020, 2022); Giblin and Mertens (2014), and magnetohydrodynamic (MHD) turbulence Caprini et al. (2009); Kahniashvili et al. (2008, 2010); Caprini et al. (2016); Wansleben et al. (2019) generate GW radiations. The resulting GW spectrum encodes information about the properties of the phase transition, such as its strength, duration, and characteristic energy scale. Observing these GWs with current and future detectors provides a direct probe of the underlying particle physics responsible for the FOPT Athron et al. (2024); Caldwell et al. (2022); Bian et al. (2021), including extensions to the Standard Model, such as additional scalar fields or symmetry-breaking mechanisms.
@@
In the early universe, FOPT occurs when the cosmic temperature drops below the critical temperature, triggering vacuum bubble nucleation in the false vacuum background. The subsequent expansion, collision, and percolation of these bubbles release significant latent heat and can generate large density fluctuations from random generation of vacuum bubbles Liu et al. (2022). If these fluctuations exceed a critical threshold, they may collapse under their own gravity to form Primordial black holes (PBHs) Hawking et al. (1982); Crawford and Schramm (1982); Kodama et al. (1982); Johnson et al. (2012); Khlopov et al. (1998). PBH is hypothesized to form in the early universe due to large density perturbations, offer a unique window into the physics of the early cosmos. Unlike stellar-mass black holes formed from collapsing stars, PBHs can span a wide range of masses, from sub-planetary to stellar and even supermassive scales, depending on their formation mechanism and the underlying physics. They have attracted considerable attention as potential candidates for dark matter, sources of high-energy phenomena, and contributors to the gravitational wave background observed by detectors such as LIGO-Virgo Sasaki et al. (2016); Bird et al. (2016); Clesse and García-Bellido (2017); Hütsi et al. (2021); Hall et al. (2020); Franciolini et al. (2022); He et al. (2024), and the planned LISA Armano et al. (2016); Amaro-Seoane et al. (2017); Seoane et al. (2023), Taiji Hu and Wu (2017); Ruan et al. (2020); Wu et al. (2021) as well as TianQin mission Luo et al. (2016, 2020); Mei et al. (2021). There are currently many proposed mechanisms for the formation of PBHs, such as inflaton ultraslow-roll Carr and Lidsey (1993); Ivanov et al. (1994), the collapse of cosmic strings Hawking (1989); Caldwell and Casper (1996); Jenkins and Sakellariadou (2020); Blanco-Pillado et al. (2021), domain walls Vachaspati (2017); Ferrer et al. (2019); Gelmini et al. (2023a, b), scalar condensates Dolgov and Silk (1993); Kasai et al. (2022); Cotner and Kusenko (2017); Martin et al. (2020), or processes in a dissipative dark sector Chang et al. (2019); Flores and Kusenko (2021); Domènech et al. (2023); Chakraborty et al. (2022).
The mass spectrum and abundance of PBHs produced through FOPT depend sensitively on the dynamics of the phase transition, such as the nucleation rate, bubble wall velocity, and phase transition strength of the transition Lewicki et al. (2024). In this work, we perform numerical simulations to investigate the density perturbation in the early universe, originating from delayed vacuum tunneling. In this scenario, specific regions of the universe undergo delayed decay of the false vacuum. As the radiation energy density decreases over time while the vacuum energy remains constant, this leads to the formation of localized energy overdensities. If these overdensities exceed a critical threshold, they may collapse into PBHs. Additionally, we compare the effects of different phase transition rates and strengths on the simulation results and further simulate the GW spectra generated by bubble expansion and collisions during the FOPT. These findings provide valuable insights into the dynamics of the early universe. For convenience, we adopt the reduced Planck mass $m_{pl}=1$ and Gravitational constant $G=1$ throughout this study.
II Density perturbations FROM FOPTs
@@
We consider a phase transition involving the scalar field $\phi$, and the associated potential can be expressed aspotential Cutting et al. (2018, 2021)
$V(\phi)=\frac{1}{2}M^{2}\phi^{2}+\frac{1}{3}\delta\phi^{3}+\frac{1}{4}\lambda\phi^{4}+\Delta V,$ (1)
@@
where $\Delta V$ is a constant used to adjust the potential’s minimum value to zero. During the transition, vacuum bubbles nucleate with an exponential rate expressed as Coleman (1977); Enqvist et al. (1992)
@@
where $\Gamma_{0}$ is the initial nucleation rate, and $\beta$ is a key parameter that dictates the rate’s increase over time. After nucleation, the field evolves according to the Klein-Gordon equation in an expanding universe Child and Giblin (2012)
@@
the spatially averaged energy density and pressure can be split into four componentsare Adshead et al. (2020)
@@
where $\rho_{K}=\frac{\phi^{\prime}}{2a^{2}}$, $\rho_{G}=\frac{(\partial_{i}\phi)^{2}}{2a^{2}}$, $\rho_{V}$, and $\rho_{r}$ are the spatially averaged kinetic energy density, gradient energy density, potential energy density, and energy density of background radiation, respectively. The PT strength is defined as $\alpha=\Delta V/\rho_{r}$, which represents the ratio of vacuum energy released to the background radiation energy density. The continuity equation can be written as
$\frac{{\mathrm{d}(\rho_{K}+\rho_{G}+\rho_{r})}}{\mathrm{d}t}+3\mathcal{H}(2\rho_{K}+\frac{2}{3}\rho_{G}+\frac{4}{3}\rho_{r})=-\frac{{\mathrm{d}\rho_{V}}}{\mathrm{d}t}\;.$ (6)
Where, $\mathcal{H}=a^{\prime}/a$ is the conformal Hubble parameter related to the usual Hubble parameter via $\mathcal{H}=aH$ with $a^{\prime}\equiv\partial a/\partial t$. The corresponding Friedmann’s equation is given as
@@
During the PT, asynchronous vacuum tunneling results in regions within each Hubble-sized volume where the false vacuum decay is delayed. In an expanding universe, while the radiation energy density decreases, the vacuum energy remains constant. This delay in false vacuum decay increases the local energy density in these regions. During the mid-to-late stages of a FOPT, density perturbations on the Hubble scale cross the horizon, leading to significant fluctuations with
@@
where $\rho_{inside}$ and $\rho_{outside}$ represent the energy densities inside and outside the delayed decay region, respectively.
The power spectrum of these perturbations is crucial for understanding the characteristics of density fluctuations. The dimensionless power spectrum can be obtained by directly differentiating the variance at the Hubble scale
$\mathcal{P_{\delta}}(k,t)=\frac{\partial\rho_{\delta}}{\partial lnk},$ (9)
where $\rho_{\delta}=(\frac{\rho-\bar{\rho}}{\bar{\rho}})^{2}$ , $\delta_{H}$ is the standard deviation of $\rho_{\delta}=(\frac{\rho-\bar{\rho}}{\bar{\rho}})^{2}$ at the Hubble horizon scale. It can be expressed as
$\delta_{H}^{2}(k,t)=\int\frac{dk}{k}\mathcal{P_{\delta}}(k,t)\;.$ (10)
The curvature perturbation power specturm can be obtained from the following equation Josan et al. (2009); Ando et al. (2018)
$\mathcal{P}_{\delta}(k,t)=\frac{16}{3}\left(\frac{k}{aH}\right)^{2}j_{1}^{2}(k/\sqrt{3}aH)\mathcal{P}_{\mathcal{R}}(k,t),$ (11)
where $j_{1}$ is a spherical Bessel function.
III GW power specturm definition
@@
GWs only require considering the transverse and traceless components. Therefore, the energy-momentum tensor for GWs is Hindmarsh et al. (2015)
@@
$\mathcal{P}_{gw}(t)=\frac{1}{24\pi^{2}\mathcal{H}^{2}V}\sum_{i,j}\int\mathrm{d}\Omega\left|\mathbf{k}\right|^{3}\left|h_{ij}^{\prime}(k)\right|^{2}$ (13)
where $V$ is the simulation volume. What is important in our simulation is that the momentum $k$ has a minimum infrared cut-off because of the discrete lattice.
IV NUMERICAL RESULTS
@@
In this simulation, we consider 64 Hubble volumes and define the last 16 to form a bubble center as postponed Hubble volumes. We calculate the energy fluctuation of each Hubble volume whose decay of false vacuum is postponed. In a rapid phase transition process, every Hubble volume can contain at least one bubble center in the end. Hence, we artificially choose that the Hubble volume getting a bubble after the time that $1/4$(maybe $1/e$ is better) of all Hubble volumes have not contained any bubble center is called the postponed Hubble volume. In a slow phase transition, due to the slow bubble nucleation rate. In this case, we can simply define the postponed Hubble volume is the region that do not contain any bubble center. Typically, in this simulation, we employ a lattice of $N^{3}=512^{3}dx$ to model a region with $LH=4$, and $H$ being the Hubble parameter. Corresponding to a total of $4^{3}$ Hubble volumes. The initial bubble radius and wall thickness are set to $12dx$ and $4dx$, respectively. The mean bubble separation is obtained as $R^{*}=(L^{3}/N_{b})^{\frac{1}{3}}$ with $N_{b}$ being the number of the generated bubbles during the FOPT processes. For the vacuum bubble nucleation probability given by Eq. (2), we explore multiple values of $\alpha=0.5,1,5,10$ and $\beta/H=6,8,10,12$. The spatial and temporal lattice spacings are defined as $\mathrm{d}x=L/N$ and $\mathrm{d}t=\mathrm{d}x/5$, respectively. The numerical calculations is are conducted with a code developed based on pystella Weiner (2021).
@@
Figure 1: EOS evolution for weak phase transition strength (left, solid line $\alpha$=0.5, dashed line $\alpha$=1) and strong phase transition strength (right, solid line $\alpha$=5, dashed line $\alpha$=10). Four colors represent different $\beta/H$ values. In all cases, $\omega$ transitions from $\omega\to-1$(vacuum energy domination) to a radiation-dominated state with $\omega=1/3$. A smaller $\alpha$ results in a longer transition time to reach $\omega=1/3$.
@@
The evolution of the average energy density of each component follows the evolution described by the continuity equation Eq. (6) and the Friedmann equation Eq. (7). As the phase transition progresses, the vacuum bubbles expand, converting the vacuum energy $\rho_{V}$ into the bubble wall energy $\rho_{w}$ (including the kinetic and gradient energy of the wall) and the background radiation energy $\rho_{r}$. Initially, bubble wall energy increases, reaching a peak before decreasing, which corresponds to changes in the phase transition rate. When the PT rate reaches its maximum, numerous vacuum bubbles form in space, with the earliest bubbles fully expanded. Consequently, the wall energy reaches its peak. As bubbles collide and merge, the bubble wall energy gradually converts into background radiation energy, eventually approaching zero. Since $\rho_{r}\propto a(t)^{-4}$, the background radiation energy density also tends towards zero after the phase transition is complete. Subsequently, the Equation of State (EoS), which is computed by $\omega=P/\rho_{tot}$, is shown in Fig. 1. It can be seen that in the beginning, the universe was dominated by vacuum energy, also known as dark energy, with a value of $\omega\to-1$. As the vacuum energy is continuously converted into the radiation energy, $\omega$ continues to increase. After the phase transition of the entire space is completed, the universe enters a radiation-dominated stage, and ultimately $\omega=1/3$.
Figure 2: Evolution diagram of $\delta_{H}(t)=\int\frac{dk}{k}\mathcal{P_{\delta}}(t)$ during FOPT process. The panels are arranged in a $2\times 2$ grid: $\alpha=0.5$ (top-left), $\alpha=1$ (top-right), $\alpha=5$ (bottom-left), and $\alpha=10$ (bottom-right). The contribution from smaller scales ($kR^{*}>1.3$) has been droped.
@@
For the density perturbation, we compute the value of $\delta_{H}$ using Eq. (10), and the corresponding evolution is presented in Fig. 2. Where, the $\delta_{H}$ for PTs of $\alpha=0.5,1$ tends to reach the maximum when $t/R^{*}\sim 1$ mostly come from moving forward of the bubble walls. Differently, the $\delta_{H}$ reach maximum at around $t/R^{*}\sim 0.5$ for strong PTs scenarios with $\alpha=5,10$ sourced from the delay of vacuum decay in different Hubble volumes, and the second peak comes from the energy transformation in the simulation volume due to the vacuum bubble walls’ moving forward behavior. Therefore, we present $\delta_{H}$ as a function of $\alpha$ for different $\beta/H$ at the two peak moments $t/R^{*}=0.5,1.2$ and plot it in Fig. 3. The left plot demonstrates that the $\delta_{H}$ increases as the PT strength increases. While, the right plot show that generally the $\delta_{H}$ increase before $\alpha\leq 1$ and decrease after $\alpha\geq 1$.
Figure 3: For $\beta/H=6,8,10,12$, the left panel shows $\delta_{H}$ at the first peak $t/R^{*}=0.5$, while the right panel presents $\delta_{H}$ at the second peak $t/R^{*}=1.2$.
@@
Figure 4: The power spectra of density perturbations for different $\beta/H$ values during the mid-stage of the phase transition. The $\alpha$ from top-left to bottom-right are 0.5, 1, 5, and 10, respectively.
Fig. 4 shows the power spectrum of density perturbations at the completion of the phase transition. For $log_{10}[kR^{*}]\textless 0.5$, the slope follows $k^{3}$, while for $0.8\textless log_{10}[kR^{*}]\textless 1.3$, the slope is $k^{-1.5}$. Notably, no clear correlation is observed between the power spectrum and the parameters $\beta/H$ or $\alpha$, suggesting that the density perturbation spectrum primarily depends on the intrinsic dynamics of the phase transition rather than its specific strength or nucleation rate. We use a function $f_{\delta}$ to fit $P_{\delta}=f_{\delta}k^{3,-1.5}$ for infra-red (IR) and ultra-violet (UV) wavenumber regions
$f_{\delta}(\alpha,\beta/H)=A(\beta/H)^{n}\alpha^{m}\;.$ (14)
For the cases where $\alpha<1$, the fitting results are $A=2.24$, $n=-0.81$, and $m=0.35$. And, for the eight cases where $\alpha>1$, the fitting results are $A=0.14$, $n=0.62$, and $m=-0.39$.
@@
Figure 5: The energy density spectra of GWs for different $\beta/H$ values at the late stage of the phase transition. The $\alpha$ from top-left to bottom-right are 0.5, 1, 5, and 10, respectively.

### v2

@@ Abstract
We conducted three-dimensional lattice simulations to study the density perturbation and gravitational waves (GWs) during first-order phase transition (FOPT). We find that for phase transition strength $\alpha>1$, the forward motion of bubble walls becomes the primary source, whereas for $\alpha<1$, the dominant contribution to the density perturbation comes from the delay of vacuum decay. Additionally, the power spectrum of density perturbations generated by the phase transition exhibits a slope of $k^{3}$ at small wavenumbers and $k^{-1.5}$ at large wavenumbers. Furthermore, we calculated the GW power spectra, which exhibit the slope of $k^{3}$ at small wavenumbers and $k^{-2}$ at large wavenumbers. Our numerical simulations confirm that slow PTs can produce PBHs and provide predictions for the GW power spectrum, offering theoretical support for GW detection.
@@ I Introduction
Primordial black holes (PBHs) are hypothesized to form in the early universe due to large density perturbations, offering a unique window into the physics of the early cosmos. There are currently many proposed mechanisms for the formation of PBHs, such as: First-order phase transition (FOPT) [Hawking:1982ga, Crawford:1982yz, Kodama:1982sf, Johnson:2011wt, Khlopov:1998nm], inflaton ultraslow-roll [Carr:1993aq, Ivanov:1994pa], the collapse of cosmic strings [Hawking:1987bn, Caldwell:1995fu, Jenkins:2020ctp, Blanco-Pillado:2021klh], domain walls [Vachaspati:2017hjw, Ferrer:2018uiu, Gelmini:2023uiu, Gelmini:2023aab, Gouttenoire:2023ftk, Gouttenoire:2023gbn], scalar condensates [Dolgov:1992pu, Kasai:2002pu, Cotner:2016cvr, Martin:2020tvv], or processes in a dissipative dark sector [Chang:2019tvv, Flores:2020drq, Domenech:2023afs, Chakraborty:2022mwu]. Unlike stellar-mass black holes formed from collapsing stars, PBHs can span a wide range of masses, from sub-planetary to stellar and even supermassive scales, depending on their formation mechanism and the underlying physics. They have attracted considerable attention as potential candidates for dark matter, sources of high-energy phenomena, and contributors to the gravitational wave background observed by detectors such as LIGO-Virgo [Sasaki:2016jop, Bird:2016dcv, Clesse:2016vqa, Hutsi:2020sol, Hall:2020daa, Franciolini:2021tla, He:2023yvl], and the planned LISA [Armano:2016bkm, LISA:2017pwj, LISA:2022yao], Taiji [Hu:2017mde, Ruan:2018tsw, TaijiScientific:2021qgx] as well as TianQin mission [TianQin:2015yph, Luo:2020bls, TianQin:2020hid].
In the early universe, FOPT proceeds through the nucleation, expansion, collision, and percolation of vacuum bubbles, and these bubbles release significant latent heat and can generate large density fluctuations from the random generation of vacuum bubbles [Liu:2021svg]. This topic has aroused sustained and growing interest because the FOPT is also an important source of stochastic gravitational wave (GW) background [Hindmarsh:2020hop]. During FOPT, the anisotropic stresses produced by vacuum bubble collisions [Kosowsky:1992rz, Kosowsky:1992vn, Kosowsky:1991ua, Child:2012qg, Cutting:2018tjt, Cutting:2020nla, Li:2023yzq, Zhao:2022cnn, Di:2020kbw], sound waves [Giblin:2013kea, Hindmarsh:2015qta, Hindmarsh:2017gnf, Hindmarsh:2013xza, Cutting:2019zws, Cutting_2022, Giblin:2014qia], and magnetohydrodynamic turbulence (MHD) [Kosowsky:2001xp, Kamionkowski:1993fg, Neronov:2020qrl, RoperPol:2019wvy, Caprini:2009yp, Kahniashvili:2008pe, Kahniashvili:2009mf, Caprini:2015zlo, Wansleben_2019] generate GW radiations. The resulting GW spectrum encodes information about the properties of the phase transition, such as its strength, duration, and characteristic energy scale. Observing these GWs with current and future detectors provides a direct probe of the underlying particle physics responsible for the FOPT [Athron:2023xlk, Caldwell:2022qsj, Bian:2021ini], including extensions to the Standard Model, such as additional scalar fields or symmetry-breaking mechanisms.
In this Letter, we perform numerical simulations to investigate the density perturbation in the early universe, originating from the FOPT. We compare the effects of different phase transition rates and strengths on the density perturbation and the GW spectra generated by bubble collisions during the FOPT. Our findings indicate that the randomness of quantum tunneling during phase transitions can induce curvature perturbations, which can be orders of magnitude larger than primordial perturbations from inflation [Liu:2021svg]. Furthermore, smaller values of $\beta/H$ favor PBH formation. In our numerical simulations, the case with $\beta/H=6$ successfully collapse into PBHs. For convenience, we adopt speed of light $c=8\pi G=1$ throughout this study.
II PBHs FROM PTs
@@
We consider a phase transition involving the scalar field $\phi$, and the associated potential can be expressed as [Cutting:2018tjt, Cutting:2020nla]
$V(\phi)=\frac{1}{2}M^{2}\phi^{2}+\frac{1}{3}\delta\phi^{3}+\frac{1}{4}\lambda\phi^{4}\;.$ (1)
@@
During the transition, vacuum bubbles nucleate with an exponential rate expressed as [Coleman:1977py, Enqvist:1991xw]
@@
where $\Gamma_{0}$ is the initial nucleation rate, and $\beta$ is a key parameter that governs the rate’s increase over time. After nucleation, the field evolves according to the Klein-Gordon equation in an expanding universe [Child:2012qg]
@@
the spatially averaged energy density and pressure can be decomposed into four components [Adshead:2019lbr]
@@
where $\rho_{K}=\frac{\phi^{\prime 2}}{2a^{2}}$, $\rho_{G}=\frac{(\partial_{i}\phi)^{2}}{2a^{2}}$, $\rho_{V}$, and $\rho_{r}$ represent the spatially averaged kinetic energy density, gradient energy density, potential energy density, and energy density of background radiation, respectively. The PT strength is defined as $\alpha=\Delta V/\rho_{r}$, which represents the ratio of vacuum energy released to the background radiation energy density. The continuity equation can be written as
$\frac{{\mathrm{d}(\rho_{K}+\rho_{G}+\rho_{r})}}{\mathrm{d}\tau}+3\mathcal{H}(2\rho_{K}+\frac{2}{3}\rho_{G}+\frac{4}{3}\rho_{r})=-\frac{{\mathrm{d}\rho_{V}}}{\mathrm{d}\tau}\;,$ (6)
where, $\mathcal{H}=a^{\prime}/a$ is the conformal Hubble parameter related to the usual Hubble parameter via $\mathcal{H}=aH$ with $a^{\prime}\equiv\partial a/\partial\tau$, the conformal time $\tau$ is related to the cosmic time $t$ via $d\tau=dt/a$. The corresponding Friedmann’s equation is given as
@@
During the PT, asynchronous vacuum tunneling results in regions within each Hubble-sized volume where the false vacuum decay is delayed. In an expanding universe, while the radiation energy density decreases, the vacuum energy remains constant. This delay in false vacuum decay increases the local energy density in these regions. During the intermediate to late stages of a FOPT, density perturbations on the Hubble scale cross the horizon, leading to significant fluctuations with
@@
where $\rho_{inside}$ and $\rho_{outside}$ represent the energy densities inside and outside the delayed decay region, respectively. If this overdensity exceeds the threshold $\delta_{c}$, the Hubble patch collapses into a PBH [Musco:2004ak, Harada:2013epa]. The numerical simulations suggest that the value of threshold is within the range of 0.40 to 0.67 [Musco:2018rwt, Escriva:2019phb], in this work, we select commonly used $\delta_{c}=0.45$ [Jedamzik:1999am, Green:2004wb, Musco:2004ak, Musco:2012au, Harada:2015yda]. When the energy density in a Hubble volume exceeds this critical threshold, the entire mass within the Hubble horizon collapses to form a PBH, and its mass can be expressed as [Liu:2021svg, Goncalves:2024vkj]
$M_{\text{PBH}}\approx\gamma V_{H}\rho_{c}=4\pi\gamma H^{-1}(t_{\text{PBH}}),$ (9)
where $\rho_{c}=3H^{2}(t_{\text{PBH}})$ is the critical density, $\gamma\leq 1$ is a correction factor associated with the specific physical processes of gravitational collapse, $t_{\text{PBH}}$ is the PBH formation time, the Hubble volume is $V_{H}=\frac{4}{3}\pi H(t)^{-3}$, and the fraction of PBHs from FOPT is [Hashino:2021qoq]
$f_{\mathrm{PBH}}=\left(\frac{H(t_{\mathrm{PBH}})}{H(t_{0})}\right)^{2}\left(\frac{a(t_{\mathrm{PBH}})}{a(t_{0})}\right)^{3}P(t_{n})\frac{1}{\Omega_{\mathrm{DM}}},$ (10)
where $\Omega_{\mathrm{DM}}=0.245$ denotes the normalized ratio of the present dark matter density to the total energy density, $t_{0}$ is the time of the first bubble nucleation, $t_{n}$ is the time for bubbles starting to nucleate in those Hubble volumes with postponed vacuum decay, and $P(t_{n})$ represents the probability that a Hubble volume collapses into a PBH at time $t_{n}$, with the following expression [Hashino:2021qoq, Liu:2021svg]
$P(t_{n})=\exp\left[-\frac{4\pi}{3}\int_{t_{0}}^{t_{n}}\frac{a^{3}(t)}{a^{3}(t_{\text{PBH}})}H^{-3}(t_{\text{PBH})}\Gamma(t)dt\right].$ (11)
Next, we define the variance of density perturbations as
$\sigma^{2}_{\delta}(k,t)=\langle(\delta(k,t)-\langle\delta(k,t)\rangle)^{2}\rangle.$ (12)
The power spectrum of these perturbations plays a crucial role in understanding the characteristics of density fluctuations. The dimensionless power spectrum can be obtained by directly differentiating the variance
$\mathcal{P_{\delta}}(k,t)=\frac{\partial\sigma^{2}_{\delta}(k,t)}{\partial lnk}.$ (13)
The relationship between the density perturbation power spectrum and the curvature perturbation power spectrum is given by [Josan:2009qn, Ando:2018qdb]
$\mathcal{P}_{\delta}(k,t)=\frac{16}{3}\left(\frac{k}{aH}\right)^{2}j_{1}^{2}(k/\sqrt{3}aH)\mathcal{P}_{\mathcal{R}}(k),$ (14)
where $j_{1}$ is a spherical Bessel function, whose value at horizon ($k_{H}=aH$, $\tau_{H}$ is the time of entry into the Hubble horizon) crossing simply reads
$\mathcal{P}_{\mathcal{R}}(k_{H})=\frac{3\mathcal{P}_{\delta}(k_{H},\tau_{H})}{16j^{2}_{1}(1/\sqrt{3})}.$ (15)
@@
GWs only require considering the transverse and traceless components. Therefore, the energy-momentum tensor for GWs is [Hindmarsh:2015qta]
@@
$\Omega_{gw}(t)=\frac{1}{24\pi^{2}\mathcal{H}^{2}V}\sum_{i,j}\int\mathrm{d}\Omega\left|\mathbf{k}\right|^{3}\left|h_{ij}^{\prime}(k,t)\right|^{2},$ (17)
where $V$ is the simulation volume.
III Numerical results
@@
In this simulation, we consider 64 Hubble volumes and define the last 16 to form a bubble center as postponed Hubble volumes. We calculate the energy fluctuation of each Hubble volume whose decay of false vacuum is postponed. In a rapid phase transition process, every Hubble volume can contain at least one bubble center in the end. Hence, we artificially choose that the Hubble volume getting a bubble after the time that $1/4$ (maybe $1/e$ is better) of all Hubble volumes have not contained any bubble center is called the postponed Hubble volume. In a slow phase transition, due to the slow bubble nucleation rate. We can simply define the postponed Hubble volume as the region that does not contain any bubble center. Typically, in this simulation, we employ a lattice of $N^{3}=512^{3}dx$ to model a region with $LH=4$. Corresponding to a total of $4^{3}$ Hubble volumes. The initial bubble radius and wall thickness are set to $12dx$ and $4dx$, respectively. The mean bubble separation is obtained as $R_{*}=(L^{3}/N_{b})^{\frac{1}{3}}$ with $N_{b}$ being the number of the generated bubbles during the FOPT processes. For the potential parameters, we choose $M=0.002$, $\delta=-0.04$, and $\lambda=7$ in order that the successful completion of the first-order PT. We simulate the quantum tunneling process in the false vacuum through stochastic bubble nucleation. For the vacuum bubble nucleation probability given by Eq. (2), we explore multiple values of $\alpha=$ 0.5, 1, 5, 10 and $\beta/H$ 6, 8, 10, 12. The spatial and temporal lattice spacings are defined as $\mathrm{d}x=L/N$ and $\mathrm{d}t=\mathrm{d}x/5$, respectively. At the end simulation, the bubble wall thickness falls within the range of $1.2dx$ to $1.6dx$, with the density fluctuation and GW spectra contributions from smaller scales ($\log_{10}[kR_{*}]>1.3$) have been dropped. The numerical calculations are conducted with a code developed based on pystella [Weiner:2021zj].
@@
Figure 1: Equation of state (EOS) evolution for weak phase transition strength (top, solid line $\alpha$=0.5, dash-dotted line $\alpha$=1) and strong phase transition strength (bottom, solid line $\alpha$=5, dash-dotted line $\alpha$=10). Four colors represent different $\beta/H$ values. In all cases, $\omega$ transitions from $\omega\to-1$ (vacuum energy domination) to $\omega=1/3$. A smaller $\alpha$ results in a longer transition time to reach $\omega=1/3$.
@@
The evolution of the average energy density of each component follows the evolution described by continuity equation Eq. (6) and the Friedmann equation Eq. (7). As the phase transition progresses, the vacuum bubbles expand, converting the vacuum energy $\rho_{V}$ into the bubble wall energy $\rho_{w}$ (including the kinetic and gradient energy of the wall) and the background radiation energy. At the end of the PT, one have the background radiation energy density dominate the energy density, see Fig. 3 in Supplemental material for details. The Equation of State (EoS), which is computed by $\omega=P/\rho_{tot}$, is shown in Fig. 1. It can be seen that in the beginning, the universe was dominated by vacuum energy, also known as dark energy, with a value of $\omega\to-1$. As the vacuum energy is continuously converted into kinetic and gradient energies, $\omega$ continues to increase. After the phase transition of the entire space is completed, the universe ultimately admit $\omega=1/3$.
Figure 2: The PBH abundance variation with respect to the PT strength parameter $\alpha$ and the inverse duration $\beta/H$.
By applying Eqs. (10) and (11), here, $H$ and $a$ are dynamically evolving parameters during the simulation process. We calculated the PBH abundance for different values of $\alpha$ and $\beta/H$, and the results are presented in Fig. 2. On the whole, we find that the PBH abundance increases with $\alpha$ and decreases with $\beta/H$, which is consistent with the conclusions of Refs. [Liu:2021svg, Kawana:2022olo, Cai:2024nln]. Notably, this dependence is stronger on $\beta/H$ and weaker on $\alpha$. In particular, when $\alpha\gtrsim 10$, the influence of $\alpha$ on the PBH abundance is further suppressed. Furthermore, the results indicate that PBH formation occurs for slow phase transitions with $\beta/H=6$, see APP. A for more details during the PT processes. This is consistent with the conclusion of Refs. [Gouttenoire:2023naa, Gouttenoire:2023bqy, Gouttenoire:2023pxh] that PBHs are more likely to form when $\beta/H<6$.
Figure 3: Evolution of $\sigma_{\delta}$ for weak phase transition strength (top, solid line $\alpha$=0.5, dash-dotted line $\alpha$=1) and strong phase transition strength (bottom, solid line $\alpha$=5, dash-dotted line $\alpha$=10).
Figure 4: Variation of the standard deviation $\sigma_{\delta}$ of accumulated overdensity perturbations with respect to the PT parameters $\alpha$ and $\beta/H$.
@@
For the density perturbation, we compute the value of $\sigma_{\delta}$ using Eq. (12), and the corresponding evolution is presented in Fig. 3. Where, the $\sigma_{\delta}$ for PTs of $\alpha=0.5,1$ tends to reach the maximum when $t/R_{*}\sim 1.2$ mostly come from the energy transformation in the simulation volume, which is influenced by the forward motion of bubble walls. Differently, the $\sigma_{\delta}$ reach maximum at around $t/R_{*}\sim 0.5$ for strong PTs scenarios with $\alpha$=5, 10 sourced from the delay of vacuum decay in different Hubble volumes. This observation is reflected in Fig. 2 and Fig. 3 of the Supplemental material, where we present details on the PT process and the energy density evolution. We plot the maximum values of $\sigma_{\delta}$ as a function of $\alpha$ for different $\beta/H$ in Fig. 4, which indicates that the magnitude of $\sigma_{\delta}$ grows as the PT strength increases.
@@
Figure 5: The power spectra of density perturbations for different $\beta/H$ values during the late stage of the phase transition. Top panel: solid line $\alpha$=0.5, dash-dotted line $\alpha$=1; Bottom panel: solid line $\alpha$=5, dash-dotted line $\alpha$=10.
Fig. 5 shows the power spectrum of density perturbations at the completion of the phase transition. For $log_{10}[kR_{*}]\textless 0.5$, the slope follows $k^{3}$, while for $0.8\textless log_{10}[kR_{*}]\textless 1.3$, the slope is $k^{-1.5}$. Notably, no clear correlation is observed between the power spectrum slope and the parameters $\beta/H$ or $\alpha$. When the curvature perturbation at the Hubble horizon scale can be derived from Eq. (15), which can be quantitatively characterized by the following fitting equation [Liu:2022lvz]
$\mathcal{P}_{\mathcal{R}}(\alpha,\beta/H)=A\alpha^{2}[f(\beta/H)]^{2}(kR_{H})^{3},$ (18)
where the $f(\beta/H)\equiv\sigma_{\delta}/\alpha$. The fitting result is $A=6.10$, compared to the value 34.5 in Ref. [Liu:2022lvz] for $\alpha\lesssim 1$ scenario.
@@
Figure 6: The energy density spectra of GWs for different $\beta/H$ values at the late stage of the phase transition. Top panel: solid line $\alpha$=0.5, dash-dotted line $\alpha$=1; Bottom panel: solid line $\alpha$=5, dash-dotted line $\alpha$=10.